OpenStudy (sette):

what is the product of (3x^2/x+2)/(5x^2-20/30x^3) ? simplify completely. Algebra 1 Honors.

OpenStudy (anonymous):

\[\frac{3x^2}{x+2} * \frac{5x^2-20}{30x^3}\]

OpenStudy (anonymous):

right?

OpenStudy (sette):

yes thats the equation.

OpenStudy (anonymous):

5x^2 - 20 can be factored as 5(x^2-4) = 5(x+2)(x-2)

razor99 (razor99):

thats the answer

OpenStudy (anonymous):

(x+2) can be cancelled out and 3x^2 can also be cancelled out so you are left with \\[\frac{5x-10}{10x}\]

OpenStudy (sette):

3x^2 can be canceled because you simplified it with 30x^3?

OpenStudy (anonymous):

yes

OpenStudy (sette):

oh okay Thank YOu .

razor99 (razor99):

hey sette check my answer

OpenStudy (radar):

\[5(x-2)\over10x\]\[x-2\over2x\]

OpenStudy (sette):

razor i dont know how you got that .

OpenStudy (radar):

I am assuming moneybird was correct and further simplified the final answer.

OpenStudy (radar):

As shown I removed the common factor and placed it outside parenthesis. then divided numerator and denominator by the 5. getting the final answer

OpenStudy (sette):

Yes Thank You. He is correct and your simplification was correct.

OpenStudy (radar):

You're welcome and good luck with these.

OpenStudy (anonymous):

(3x^2/x+2)/(5x^2-20/30x^3) (3x^2/x+2) * ( 30x^3/ (5x^2-20) = 90 x^3/ 5 ( x^3 - 8) = 18 x^3/ ( x^3 - 8)